OpenStudy (anonymous):

During the first part of a trip,a canoeist travels 71 miles at a certain speed. The conoeist travels 10 miles on the second part of the trip at a speed of 5mph slower. The total time for the trip is 5 hours. What was the speed of each part of the trip.

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

ok so distance1=71 speed1=x distance2=10 speed2=x-5 since speed=distance/time we can write it like this time=distance/speed following me?

OpenStudy (anonymous):

since total time is 50 we can write equation as 50=(71/x)+(10/x-5)

OpenStudy (anonymous):

ok so am I suppse to solve or just write the equation?

OpenStudy (anonymous):

ill solve it for u dont worry but do u understand it? lemme know if u wanna solve it urself

OpenStudy (anonymous):

I think I understand. No go ahead.

OpenStudy (anonymous):

gimme a minute i think i made a calc error trying to go through what i go wrong

OpenStudy (anonymous):

is tht right?

OpenStudy (anonymous):

mathtard u there?

OpenStudy (anonymous):

it says wrong

OpenStudy (anonymous):

noooooooo ok ill check calc again

OpenStudy (anonymous):

k

OpenStudy (anonymous):

sorry if its not 5.27

OpenStudy (anonymous):

then i dont know

OpenStudy (anonymous):

5.27 and 0.27 make sense

OpenStudy (anonymous):

sorry once again if i wasnt helpful

OpenStudy (aravindg):

hahd can u help me?/

OpenStudy (anonymous):

hey did u ditch me

OpenStudy (anonymous):

no im here

OpenStudy (anonymous):

is the answer right?

OpenStudy (anonymous):

5.27

OpenStudy (anonymous):

and 0.27

OpenStudy (aravindg):

helo me in doing physics

OpenStudy (anonymous):

ok aravind post question

OpenStudy (anonymous):

17.03 and 5.27

OpenStudy (anonymous):

tht doesnt make sense

OpenStudy (anonymous):

if he was going at 17.03 then u subtract 5 u should get 12.03

OpenStudy (saifoo.khan):

Still Stuck?

OpenStudy (anonymous):

i got 5.27 saifoo

OpenStudy (anonymous):

then subtracted 5 to get 0.27 i dont know how u would get 17.03

OpenStudy (anonymous):

saifoo u there u?

OpenStudy (anonymous):

got ditched again

OpenStudy (anonymous):

sorry didn't mean to ditch you.